OpenStudy (anonymous):

Factor completely: 5x(x + 3) + 6(x + 3) Can some one please help. i dont understand it at all

hero (hero):

Hint: ba + ca = a(b+c) In this case, a = (x + 3)

OpenStudy (anonymous):

eq. is 5x^2+15x+6x+18

hero (hero):

@deadman340 , we're trying to factor completely here.

hero (hero):

Similar Example: 1. Recognize that we have two terms and x + 7 is the common expression: 2x(x + 7) + 3(x+7) 2. Factor out x + 7 by way of distributive rule: a = (x+7) b = 2x c = 3 Since ab + ac = a(b+c) Therefore: 2x(x+7) + 3(x+7) = (x+7)(2x + 3) 3. That's it Okay, your turn.